OpenStudy (anonymous):

Describe the graphs of the functions f(x) = 2x + 1 and g(x) = -2x + 1. Compare and contrast the domain and range of f(x) and g(x).

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(x) }\) and \(\color{#000000 }{ \displaystyle -f(x) }\) are reflections of each other across the x-axis.

OpenStudy (solomonzelman):

Therefore, \(\color{#000000 }{ \displaystyle f(x)=2x }\) \(\color{#000000 }{ \displaystyle g(x)=-2x }\) Are reflections of each other across the x-axis.

OpenStudy (solomonzelman):

And so would be \(\color{#000000 }{ \displaystyle f(x)=2x+1 }\) \(\color{#000000 }{ \displaystyle g(x)=-2x-1 }\) reflections of each other across the x-axis.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle g(x)=-2x+1=(-2x-1)+2 }\)

OpenStudy (solomonzelman):

So you can tell that \(\color{#000000 }{ \displaystyle g(x) }\) is a reflection across the x-axis of the \(\color{#000000 }{ \displaystyle f(x) }\), and then \(\color{#000000 }{ \displaystyle g(x) }\) is shifted by .... (you should tell me this)

OpenStudy (anonymous):

f(x)?

OpenStudy (anonymous):

@SolomonZelman